Project

General

Profile

Wiki » History » Version 29

Version 28 (Ken Lutz, 08/01/2016 11:14 AM) → Version 29/84 (Ken Lutz, 08/01/2016 04:41 PM)

# GDP Deployment Notes

## Data users email list created
If you would like to be notified of data anomalies or changes in the sensor actuator send an email to lutz@berkeley.edu asking to be added to the email list, swarm-gdp-data-users@lists.eecs.berkeley.edu

**log naming convention:** Unless otherwise noted, the name of the corresponding logs are derived from the physical location (currently `bwrc` or `swarmlab`) and the lowercase MAC address:

edu.berkeley.eecs.<location>.device.<mac-addr>

For example, a sensor with MAC address `C0:98:E5:30:00:03` in the immersion room (which is in swarmlab) will be named `edu.berkeley.eecs.swarmlab.device.c098e5300003`

## BWRC Server Room
**Sensors and Gateways**
BBB - UHKBBB004.eecs.berkeley.edu 128.32.63.176
Rear Wall BLEE - C0:98:E5:30:00:36
Mid Wall BLEE - C0:98:E5:30:00:5D
Hot Aisle BLEE- C0:98:E5:30:00:54
BLINK - C0:98:E5:90:00:1E
Lab Rear Wall BLEE - C0:98:E5:30:00:09
Power Blade - Utility power to sump pump - C0:98:E5:70:00:8F
Power Blade - Utility power to Air Compressor - C0:98:E5:70:90
USB Microphone - mounted on BBB

## BWRC Office Area
**Sensors and Gateways**
Swarmnuc1002.eecs.berkeley.edu
**BLEE**
Main seating area - C0:98:E5:30:00:A5
Conference Room , Rabaey Room - C0:98:E5:30:00:A3
Admin Office - C0:98:E5:30:00:A4
Conference Room - Brodersen room - C0:98:E5:30:00:A2
Ken's Office - C0:98:E5:30:00:A1
Brian Office above thermostat - C0:98:E5:30:00:A0
Center of student seating area, above thermostat - C0:98:E5:30:00:37
**Blink**
Brodersen Conference Room C0:98:E5:90:00:87
Rabaey Conference room - C0:98:E5:30:00:94

**Operational Notes**
6-17-16 2:59PM Repaired broken Blink in server room. Was reading all zero, now detecting movement correctly.
7/11/16 Fixed powerblade on air compressor - was installed backward. Data collected after 7/11 looks correct. Prior was in error. Note that energy accumulated goes to zero when powerblade was disconnected for inspection. Good marker.
7/14/16 9:29AM Correlated spike in temperature and humidity for a short period of time across all BLEES in the server room. Probably related to someone working in the room and having propped the door open.
7/17/16-7/19/16: Missed about 2 days worth of data due to bug in the MQTT-gateway.
7/19/16 - :BWRC BLEE, Lab Rear wall C0:98:E5:30:00:09
7/19/16 4:05 PM - 4:19PM
8/1/16 BBB004 down for maintenance 2-5pm.


Put a cover over the BLEE causing the light level to go to zero. Temp and humidity are down as well.

7/19/16 4:20PM - 4:31PM
BWRC BLEE mid wall C0:98:E5:30:00:5D

Put a cover over BLEE in server room, mid wall. Light level goes to zero, temp goes up
Other sensors in the server room do not show similar readings

7/25/16 - Added sensors and gateway to BWRC office area. New batteries

## Cory Hall Swarm Lab and Immersion Room
**Servers and Sensors**
Immersion Room
Primary projector PowerBlade ~~C0:98:E5:70:00:2B~~ C0:98:E5:70:00:8D
Secondary projector PowerBlade C0:98:E5:70:00:8B
Speaker Camera PowerBlade C0:98:E5:70:00:8E
Audience Camera PowerBlade C0:98:E5:70:00:88
BLEE, rear camera pillar C0:98:E5:30:00:03
Front wall adjacent to screen Blink C0:98:E5:90:00:0A
UHKBBB001.eecs.berkeley.edu 128.32.33.236

Chair BLEES C0:98:E5:30:0: {58, 5c, 5F, 7a, 7b, 74, 75, 76, 77, 78}

**Operational Notes**
7/12/16 2:59PM, powerblade on Speaker camera disconnected for inspection and reconnected.
Note that accumulated energy goes to zero at this time.
7/13/16 10AM, powerblade on Immersion room primary projector re-installed, now in correct orientation. Readings from 10am on look correct.
7/14/16 9:22:58AM - Introduced an anomaly by heating the Immersion room BLEE located on the rear pillar with a hot air gun for about 4 min. Temperature and humidity impacted.
7/14/16 9:46AM - Speaker camera was left on since 7/12 2:26pm. Turned off at 7/14 9:46AM
7/15/16 - 9:40AM Primary projector powerblade seems to have a loose connection. No readings from 7/14 3:23PM until re-seated at 7/15 9:40AM
7/18/16 - 11:30AM BLEES installed on conference table chairs w/ new batteries
7/17/16-7/19/16: Missed about 2 days worth of data due to bug in the MQTT-gateway.
7/19/16: Beaglebone gateway is got powered off. Last data collected was around 4:56PM. Restarted at 10:50PM
7/20/16: Rear Camera pillar BLEE - Battery replaced 3:26PM
7/26/16 Main projector power blade failed - replaced with new unit (new MAC)
8/1/16 BBB001 down for maintenance


## Jacobs Hall
Gateways
ML-100 - swarmnuc1005.eecs.berkeley.edu B8:AE:ED:74:09:3D was port 10D-120C-03F10-D room 120
BBB - uhkbbb005.eecs.berkeley.edu port 10D-332-04F17-D Room 332
BBB uhkbbb006.eecs.berkeley.edu 54:4a:16:e0:8b:a1 10D-120C-03F10-D room 120
BBB uhkbbb007.eecs.berkeley.edu 6c:ec:eb:5c:64:5b Room 10 10D-10B-02f16-D

### Room 110
**Powerblades**
c0:98:e5:70:01:13 Type A 3D printer 10026 room 110
c0:98:e5:70:01:22
c0:98:e5:70:01:23 Type A 3D printer 10047 room 110
c0:98:e5:70:01:25
c0:98:e5:70:01:26 Laser 2 room 110C
c0:98:e5:70:01:27 Type A 3D printer 10097 room 110
c0:98:e5:70:01:29 Type A 3D printer 10213 room 110
c0:98:e5:70:00:e2 Laser 1 room 110C
c0:98:e5:70:00:e5 Type A 3D printer 10035 room 110
c0:98:e5:70:00:d3 Type A 3D printer 10098 room 110
c0:98:e5:70:00:d7 Type A 3D printer 10025 room 110
c0:98:e5:70:00:f7 Type A 3D printer 10038 room 110
c0:98:e5:70:00:db - Type A 3D printer 10094 room 110
c0:98:e5:70:01:31 Type A 3D printer 10034 room 110
**
BLEE**
c0:98:e5:30:00:57 room 110 by thermostst
c0:98:e5:30:00:62 room 10C Woodshop

### Room 332

**PowerBlades**
c0:98:e5:70:00:8a Carbon 3D printer room 332
c0:98:e5:70:01:0c Dimension 2 (Bonnie) room 332
c0:98:e5:70:01:2e Dimension 1 (Clyde) room 332

**BLEE**
c0:98:e5:30:00:2f
c0:98:e5:30:00:55 room 120
c0:98:e5:30:00:50 room 332 Fortus 380 air exhaust vent
c0:98:e5:30:00:51 room 332 center of room

**Blink**
c0:98:e5:90:00:85 room 332

c0:98:e5:90:00:17
c0:98:e5:90:00:86
c0:98:e5:90:00:94
c0:98:e5:90:00:96

c0:98:e5:90:00:87

### Room 10C
**Blink**
c0:98:e5:90:00:81 - wood shop room 10C

**Operational Notes**
Sensors installed, new batteries 7/15/16